OpenStudy (anonymous):

if the volume of a shere is 10293.33 cm^3, what is the diameter? Use 3.14 to approximate n. Round answeres to the nearest hundreth.

OpenStudy (callisto):

Apply the formula \(V=\frac{4}{3} \pi r^3\). Put V=10293.33 and \(\pi = 3.14\) into the formula. \[10293.33=\frac{4}{3} (3.14) r^3\]Now, can you solve r?

OpenStudy (anonymous):

is your r 4.186?

OpenStudy (callisto):

Nope...

OpenStudy (anonymous):

so how do you solve it then ?

OpenStudy (callisto):

First, divide both sides by 4/3 Next, divide both sides by 3.14 Lastly, take cube root for both sides.

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

13.50 using Callisto's equation.

OpenStudy (anonymous):

callisto i did that and i got the wrong answer can you show me ?

OpenStudy (anonymous):

Show your steps so we can help you find out where you went wrong. Do you have an answer guide BTW?

OpenStudy (callisto):

\[10293.33=\frac{4}{3} (3.14) r^3\]\[7719.9975=(3.14) r^3\]\[2458.59793= r^3\]\[r=13.50\]

OpenStudy (callisto):

I think you should show your steps too, just as @Wired said.

OpenStudy (anonymous):

where i went wrong was i did (4/3)*10293.33 = 4/3(4/3)

OpenStudy (callisto):

Ah... Divide both sides by 4/3 is actually multiply both sides by 3/4

OpenStudy (callisto):

BTW, what you've found is the radius only!!!! Diameter = radius x 2!

OpenStudy (callisto):

@edwina3695